# Design of Bayesian Clinical Trials with Clustered Data and Multiple Endpoints

### v2

@@ 2.1 Operating Characteristics with Multiple Estimands
Given our simulation results, the probability in (3) is estimated as
$\dfrac{1}{m}\sum_{r=1}^{m}\mathbb{I}\left\{\nu(\mathcal{D}_{c,r})=1\right\},$ (4)
where $\mathcal{D}_{c,r}$ are generated using $\boldsymbol{\theta}_{r}$ obtained via $\Psi$. If $H_{1}$ is true, we want the decision criteria to be satisfied. The probability of making the correct decision based on the data is $\mathbb{E}_{\Psi_{1}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]$ where $\Psi_{1}$ is a probability model such that $H_{1}$ is true. This probability is estimated using (4) when $\mathcal{D}_{c,r}$ are generated using $\boldsymbol{\theta}_{r}$ obtained via $\Psi_{1}$. In standard contexts where $\boldsymbol{\gamma}$ is a vector of superiority thresholds, the probability of making the correct decision given a set of parameter values resulting from a degenerate $\Psi_{1}$ is referred to as power. The term assurance is commonly used if $\Psi_{1}$ incorporates uncertainty about the parametric assumptions used to generate the data (O’Hagan and Stevens, 2001). If $H_{0}$ is instead true, we do not want the decision criteria to be satisfied. The probability of making the incorrect decision is defined as $\mathbb{E}_{\Psi_{0}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]$ where $\Psi_{0}$ is a probability model such that $H_{0}$ is true. Using $\Psi_{0}$ instead of $\Psi_{1}$, this probability can be estimated as in (4). The probability of making the incorrect decision is the expected type I error rate in standard contexts where $\boldsymbol{\gamma}$ represents superiority thresholds.
The decision thresholds $\boldsymbol{\gamma}$ bound the probability of making the incorrect decision in a trial. Design methods based on simulating data allow us to select thresholds that account for the induced dependence structure across the multiple outcomes. The sample size $c$ is selected to ensure the trial has a large enough probability of making the correct decision. For every value of $c$ considered, we must obtain a collection of $\{\boldsymbol{\tau}(\mathcal{D}_{c,r})\}_{r=1}^{m}$ values via simulation to estimate this probability in (4). The process to obtain the $\{\boldsymbol{\tau}(\mathcal{D}_{c,r})\}_{r=1}^{m}$ values is often computationally intensive. However, we could reduce the computational burden by using previously estimated sampling distributions of $\boldsymbol{\tau}(\mathcal{D}_{c})$ to estimate operating characteristics at new $c$ values. We could use this process to conduct SSD for Bayesian trials with substantially fewer simulation repetitions. We propose such a method for trial design in this paper and begin its development after a brief discussion of marginal estimands in Section 2.2.
2.2 Marginal Estimands
The marginal estimands $\boldsymbol{\delta}(\boldsymbol{\theta})$ can be expressed as a contrast of population quantities for each treatment. We let $A_{i}\in\{0,1\}$ denote the treatment assignment associated with the $i^{\text{th}}$ observation, where $A_{i}=1$ and $A_{i}=0$ respectively represent being randomized to the treatment and control groups. We let $\boldsymbol{\mu}(\boldsymbol{\theta};A)$ be the marginal mean corresponding to the treatment group $A$. We represent the marginal estimands by
$\boldsymbol{\delta}(\boldsymbol{\theta})=\boldsymbol{\mu}(\boldsymbol{\theta};A=1)-\boldsymbol{\mu}(\boldsymbol{\theta};A=0),$ (5)
where we use a difference-based contrast without loss of generality. For clustered data, $Kq$ unobserved latent effects ${\bf{W}}_{c\times Kq}$ account for the dependence between outcomes from the same cluster, where there are $q$ latent effects for each of the $K$ outcomes. We assume that these latent effects are not shared between observations from different clusters. As a result, we make the standard assumption that outcomes in different clusters are independent. The joint posterior of $\boldsymbol{\theta},{\bf{W}}~{}|~{}\mathcal{D}_{c}$ generally depends on $p$ observed covariates ${\bf{X}}_{N\times p}$. Thus, $\boldsymbol{\theta}$ may contain regression coefficients corresponding to the observed covariates and distributional parameters for the latent effects.
Samples from the posterior of $\boldsymbol{\theta},{\bf{W}}~{}|~{}\mathcal{D}_{c}$ give rise to posterior samples from the conditional parameter $\boldsymbol{\mu}(\boldsymbol{\theta};A,{\bf{X}},{\bf{W}})$ with respect to fixed patterns for the covariates ${\bf{X}}$ and latent effects ${\bf{W}}$. Because the estimands in clinical trials are often non-collapsible (Daniel et al., 2021; Willard et al., 2024), we marginalize the posterior samples of $\boldsymbol{\mu}(\boldsymbol{\theta};A,{\bf{X}},{\bf{W}})$ with respect to the distributions of ${\bf{X}}$ and ${\bf{W}}$ to obtain samples from the posterior of
[…]
The procedure to implement this marginalization is based on reweighting according to repeated sampling from Dirichlet distributions; its details are described in Appendix B of the online supplement. We use this procedure to approximate the posterior distribution of $\boldsymbol{\delta}(\boldsymbol{\theta})$.
Lastly, we underscore why our methodology provides the option to marginalize over the distribution of the latent effects ${\bf{W}}$. Most cluster-specific models are predicated on the existence of latent risk groups that are indexed by the cluster-specific effects (Neuhaus et al., 1991). These latent risk groups may not be balanced for all treatment groups in trials with clustered data, impacting the validity of the trial’s conclusions. Since these cluster-specific effects are not observed, it is difficult to determine whether they are balanced between treatment groups. We therefore conduct inference by treating ${\bf{W}}$ as additional variables that we marginalize over.
3 A Proxy to the Joint Sampling Distribution
To motivate our SSD procedure proposed in Section 4, we create a proxy to the joint sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$. These proxies are needed for the theory that underpins our proposed methodology. However, our methods do not directly use these proxies and instead estimate true sampling distributions of $\boldsymbol{\tau}(\mathcal{D}_{c})$ by simulating samples $\{\mathcal{D}_{c,r}\}_{r=1}^{m}$ and approximating posterior probabilities as described in Section 2.1. Our proxies are predicated on an asymptotic approximation to the posterior of $\boldsymbol{\delta}(\boldsymbol{\theta})$ based on the Bernstein-von Mises (BvM) theorem (van der Vaart, 1998) that is detailed in Lemma 1.
We require that the four conditions for the BvM theorem are satisfied to apply our methodology. The first assumption is an independently and identically distributed assumption. We require that independence assumptions are satisfied at the cluster level but not at the observation level. The next two conditions concern the likelihood function and are weaker than the regularity conditions for the asymptotic normality of the maximum likelihood estimator (MLE) (Lehmann and Casella, 1998). For reasons described shortly, our methodology also requires that those regularity conditions are satisfied. The final condition for the BvM theorem concerns the analysis prior $p(\boldsymbol{\theta})$. This prior must be absolutely continuous with positive density in a neighbourhood of the true value for $\boldsymbol{\theta}$. For simulation repetition $r$, this true value of $\boldsymbol{\theta}_{r}\sim\Psi$ is incorporated into Lemma 1.
Lemma 1.
Assume the conditions described above are satisfied for $\boldsymbol{\theta}_{r}\sim\Psi$. Let $\hat{\boldsymbol{\delta}}^{{}_{(c)}}_{r}=\boldsymbol{\delta}(\hat{\boldsymbol{\theta}}^{{}_{(c)}}_{r})$ be the maximum likelihood estimate for $\boldsymbol{\delta}(\boldsymbol{\theta})$ expressed in terms of the number of independent clusters $c$. As $c$ increases, a large-sample approximation to the posterior of $\boldsymbol{\delta}(\boldsymbol{\theta})$ takes the form
$\mathcal{N}\left(\hat{\boldsymbol{\delta}}^{{}_{(c)}}_{r},c^{-1}\boldsymbol{\Lambda}(\boldsymbol{\theta}_{r})\right),$ (6)
where $\boldsymbol{\Lambda}(\boldsymbol{\theta}_{r})$ is related to the Fisher information matrix $\mathcal{I}(\boldsymbol{\theta})$.
In Appendix A.1 of the supplement, we prove Lemma 1 and specify the form of the matrix $\boldsymbol{\Lambda}(\boldsymbol{\theta}_{r})$. We also carefully justify that Lemma 1 holds true for marginal estimands as defined in (5) when the cluster sizes $\{n_{j}\}_{j=1}^{c}$ are drawn according to a given probability distribution. Here, we use the approximation in (6) for theoretical development. The theory in Appendix A additionally demonstrates that the approximate sampling distribution of the MLE $\hat{\boldsymbol{\delta}}^{{}_{(c)}}~{}|~{}\boldsymbol{\theta}=\boldsymbol{\theta}_{r}$ is $\mathcal{N}(\boldsymbol{\delta}_{r},c^{-1}\boldsymbol{\Lambda}(\boldsymbol{\theta}_{r}))$ under the regularity conditions in Lehmann and Casella (1998), where $\boldsymbol{\delta}_{r}=\boldsymbol{\delta}(\boldsymbol{\theta}_{r})$. For theoretical purposes only, a single realization from this $K$-dimensional multivariate normal distribution could be generated using conditional cumulative distribution function (CDF) inversion and a point $\boldsymbol{u}=\{u_{k}\}_{k=1}^{K}\in[0,1]^{K}$. We obtain the first component $\hat{\delta}^{{}_{(c)}}_{r,1}$ as the $u_{1}$-quantile of the sampling distribution of $\hat{\delta}^{{}_{(c)}}_{1}~{}|~{}\boldsymbol{\theta}_{r}$. For the remaining components, we generate $\hat{\delta}^{{}_{(c)}}_{r,k}$ as the $u_{k}$-quantile of the sampling distribution of $\hat{\delta}^{{}_{(c)}}_{k}~{}|~{}\{\hat{\delta}^{{}_{(c)}}_{s}=\hat{\delta}^{{}_{(c)}}_{r,s}\}_{s=1}^{k-1},\hskip 0.1pt\boldsymbol{\theta}_{r}$.
Implementing this process with a sequence of $m$ points $\{\boldsymbol{u}_{r}\}_{r=1}^{m}\sim\mathcal{U}([0,1]^{K})$ simulates a sample from the approximate sampling distribution of $\hat{\boldsymbol{\delta}}^{{}_{(c)}}$ according to $\Psi$. We substitute this sample $\{\hat{\boldsymbol{\delta}}^{{}_{(c)}}_{r}\}_{r=1}^{m}$ into the posterior approximation in (6) to yield samples of posterior probabilities. For each scalar estimand in $\boldsymbol{\delta}(\boldsymbol{\theta})$, we approximate the probability in the $k^{\text{th}}$ row of (2) as
$\tau^{{}_{(c)}}_{r,k}=\Phi\left(\dfrac{\delta_{U,k}-\hat{\delta}^{{}_{(c)}}_{r,k}}{\sqrt{c^{-1}\Lambda_{k,k}(\boldsymbol{\theta}_{r})}}\right)-\Phi\left(\dfrac{\delta_{L,k}-\hat{\delta}^{{}_{(c)}}_{r,k}}{\sqrt{c^{-1}\Lambda_{k,k}(\boldsymbol{\theta}_{r})}}\right),$ (7)
where $\Phi(\cdot)$ is the standard normal CDF. The collection of $\{\tau^{{}_{(c)}}_{r,k}\}_{k=1}^{K}$ values corresponding to $\{\boldsymbol{u}_{r}\}_{r=1}^{m}\sim\mathcal{U}([0,1]^{K})$ and $\{\boldsymbol{\theta}_{r}\}_{r=1}^{m}\sim\Psi$ define our proxy to the joint sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$. We acknowledge that this proxy to the sampling distribution of posterior probabilities relies on asymptotic results, and it may differ materially from the true sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ for finite $c$.
The proxy sampling distribution therefore only motivates our theoretical result in Theorem 1. This result guarantees that the logit of $\tau^{{}_{(c)}}_{r,k}$ is an approximately linear function of $c$ for all scalar estimands in $\boldsymbol{\delta}(\boldsymbol{\theta})$. We later adapt this result to assess the operating characteristics of a trial across a broad range of sample sizes by estimating the true sampling distribution of $\tau(\mathcal{D}_{c})$ at only two values of $c$. Each $\tau^{{}_{(c)}}_{r,k}$ value in the proxy sampling distribution depends on the value for $\hat{\boldsymbol{\delta}}^{{}_{(c)}}_{r}$, which depends on the sample size $c$, the parameter value $\boldsymbol{\theta}_{r}$, and the point $\boldsymbol{u}_{r}$. We emphasize that $\boldsymbol{u}_{r}\sim\mathcal{U}([0,1]^{K})$ is a stochastic conduit for the data and that $\boldsymbol{\theta}_{r}\sim\Psi$ is stochastic when $\Psi$ is nondegenerate. In Theorem 1, we fix both the $\boldsymbol{\theta}_{r}$ and $\boldsymbol{u}_{r}$ values to explore the behavior of $\tau^{{}_{(c)}}_{r,k}$ as a deterministic function of $c$.
Theorem 1.
For any $\boldsymbol{\theta}_{r}\sim\Psi$, let the conditions for Lemma 1 be satisfied. Define $\emph{logit}(x)=\emph{log}(x)-\emph{log}(1-x)$. We consider a given point $\boldsymbol{u}_{r}\in[0,1]^{K}$ and distributions for ${\bf{X}}$, ${\bf{W}}$, and the cluster size $n_{j}$. For $k=1,\dots,K$, the functions $\tau^{{}_{(c)}}_{r,k}$ in (7) are such that
  1. (a)
    $\boldsymbol{\tau}^{{}_{(c)}}_{r,k}=\Phi\left(a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})\sqrt{c}+b_{k}(\boldsymbol{u}_{r})\right)-\Phi\left(a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})\sqrt{c}+b_{k}(\boldsymbol{u}_{r})\right)$, where $a_{k}(\delta,\boldsymbol{\theta}_{r})=(\delta-\delta_{r,k})/\sqrt{\Lambda_{k,k}(\boldsymbol{\theta}_{r})}$ and $b_{k}(\cdot)$ are functions that do not depend on $c$.
  2. (b)
    $\lim\limits_{c\rightarrow\infty}\dfrac{d}{dc}~{}\emph{logit}\left(\boldsymbol{\tau}^{{}_{(c)}}_{r,k}\right)=(0.5-\mathbb{I}\{\delta_{r,k}\notin(\delta_{L,k},\delta_{U,k})\})\hskip 1.0pt\times\emph{min}\{a_{k}(\delta_{U,k},\boldsymbol{\theta}_{r})^{2},a_{k}(\delta_{L,k},\boldsymbol{\theta}_{r})^{2}\}$.
We prove parts $(a)$ and $(b)$ of Theorem 1 in Appendices A.2 and A.3. Hagar and Stevens (2025) considered a similar theoretical result that did not accommodate dependent data nor account for marginal or multiple estimands. The methods proposed in this paper therefore accommodate a broader set of data-driven comparisons in Bayesian trials. We now discuss the practical implications of Theorem 1. The limiting derivative in part $(b)$ is a constant that does not depend on $c$. For each marginal probability in the joint proxy sampling distribution, the linear approximation to $l^{{}_{(c)}}_{r,k}=\text{logit}(\tau^{{}_{(c)}}_{r,k})$ as a function of $c$ is thus a good global approximation for large sample sizes. This linear approximation should be locally suitable for a range of smaller sample sizes. Therefore, the quantiles of the sampling distribution of $l^{{}_{(c)}}_{r,k}$ change linearly as a function of $c$ when $\boldsymbol{\theta}_{r}$ is held constant across simulation repetitions. In Section 4, we exploit and adapt this linear trend in the proxy sampling distribution to flexibly model the logits of $\boldsymbol{\tau}(\mathcal{D}_{c})$ as linear functions of $c$ when independently simulating samples $\mathcal{D}_{c,r}$ according to $\boldsymbol{\theta}_{r}\sim\Psi$. While the proxy sampling distribution is predicated on asymptotic results, we illustrate the good performance of our SSD procedure with finite sample sizes $c$ in Section 5.
4 Sample Size Determination Procedure
We generalize the results from Theorem 1 to develop a procedure for Bayesian SSD that can easily be implemented when estimating posterior probabilities by simulating data $\mathcal{D}_{c}$. This procedure is described in Algorithm 1. It performs well for moderate and large sample sizes, and we estimate the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ at only two values of $c$: $c_{0}$ and $c_{1}$. This initial sample size $c_{0}$ can be selected based on the anticipated budget for the trial. In Algorithm 1, we add a subscript to $\mathcal{D}_{c,r}$ to distinguish whether the data are generated according to the model $\Psi_{0}$ or $\Psi_{1}$ defined in Section 2.1. That is, $\mathcal{D}_{c,0,r}$ and $\mathcal{D}_{c,1,r}$ represent data generated according to $\Psi_{0}$ and $\Psi_{1}$, respectively. The subscript on the number of clusters $c$ is not related to the hypothesis.
In addition to the choices discussed in Section 2, we specify a distribution for the $n_{j}\times K$ matrix of responses ${\bf{Y}}_{j}$ at the cluster level, denoted broadly by $p({\bf{Y}}_{j};\boldsymbol{\theta})$. We must also specify parameters to generate the cluster sizes $\{n_{j}\}_{j=1}^{c}$, covariates ${\bf{X}}_{N\times p}$, and latent effects ${\bf{W}}_{c\times Kq}$. We collectively denote these parameters as $\boldsymbol{\zeta}$. To implement Algorithm 1, we define criteria for the probabilities of making the correct and incorrect decisions in the trial. Under $\Psi_{1}$ where $H_{1}$ is true, we want $\mathbb{E}_{\Psi_{1}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]\geq\Gamma_{1}$. We want $\mathbb{E}_{\Psi_{0}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]\leq\Gamma_{0}$ under $\Psi_{0}$ where $H_{0}$ is true. Algorithm 1 details a general application of our methodology with a non-adaptive design, and we later describe potential modifications.
Algorithm 1 Procedure to Determine the Number of Clusters
1:procedure ClusterSSD($p({\bf{Y}}_{j};\boldsymbol{\theta})$, $\boldsymbol{\delta}(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $p(\boldsymbol{\theta})$, $\Psi_{0}$, $\Psi_{1}$, $m$, $c_{0}$, $\boldsymbol{\zeta}$, $\Gamma_{0}$, $\Gamma_{1}$)
2:     Compute $\{\boldsymbol{\tau}(\mathcal{D}_{c_{0},0,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{0}$.
3:     Choose $\boldsymbol{\gamma}$ to ensure $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\nu(\mathcal{D}_{c_{0},0,r})=1\}\leq\Gamma_{0}$.
4:     Compute $\{\boldsymbol{\tau}(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{1}$.
5:     If $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\nu(\mathcal{D}_{c_{0},1,r})=1\}\geq\Gamma_{1}$, choose $c_{1}<c_{0}$. If not, choose $c_{1}>c_{0}$.
6:     Compute $\{\boldsymbol{\tau}(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{1}$.
7:     for $k$ in 1:$K$ do
8:         Sort logits $\{l_{k}(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ and $\{l_{k}(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$ and pair their corresponding order statistics with           linear approximations to obtain $m$ $\hat{l}_{k}(\mathcal{D}_{c,1})$ estimates for new $c$ values.
9:     Obtain $\{\hat{\boldsymbol{\tau}}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ as the inverse logits of the estimates $\{\hat{l}_{k}(\mathcal{D}_{c,1})\}_{k=1}^{K}$ whose linear approximations      were defined using $\{\boldsymbol{\tau}_{k}(\mathcal{D}_{c_{0},1,r})\}_{k=1}^{K}$.
@@ 4 Sample Size Determination Procedure
10:     Find $c_{2}$, the smallest $c\in\mathbb{Z}^{+}$ such that $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\hat{\nu}(\mathcal{D}_{c,1,r})=1\}\geq\Gamma_{1}$.
@@
We now elaborate on several steps of Algorithm 1. In Line 3, we choose a suitable vector of thresholds $\boldsymbol{\gamma}$ to ensure the estimate for $\mathbb{E}_{\Psi_{0}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]$ based on the formula in (4) is at most $\Gamma_{0}$. There are many combinations of the individual thresholds $\{\gamma_{k}\}_{k=1}^{K}$ that will attain the desired bound for $\mathbb{E}_{\Psi_{0}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]$. Without loss of generality, we assume that the estimands in $\boldsymbol{\delta}(\boldsymbol{\theta})$ are indexed according to their importance (i.e., the first scalar estimand corresponds to the most important outcome). We recommend obtaining the vector $\boldsymbol{\gamma}$ by sequentially choosing $\gamma_{k}$ from $k=1$ to $K$ to satisfy criteria for the individual comparisons between $\tau_{k}(\mathcal{D}_{c})$ and $\gamma_{k}$. In Line 5, we choose a second sample size $c_{1}$ at which to estimate the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ under $\Psi_{1}$. This sample size could be selected using the limiting slopes from Theorem 1 (see Hagar and Stevens (2025) for a similar approach). An alternative strategy is to select $c_{0}$ and $c_{1}$ to facilitate visualization of the operating characteristics over a range of relevant sample sizes. We elaborate on the rationale for selecting $\boldsymbol{\gamma}$, $c_{0}$, and $c_{1}$ for our motivating example in Section 5.
@@
We emphasize that all posterior probabilities approximated in Lines 2 to 6 of Algorithm 1 are obtained by simulating data according to $\Psi_{0}$ or $\Psi_{1}$. One strength of our methodology is that it can be integrated with any computational or analytical method to estimate posterior probabilities. Lines 7 and 8 compute logits of these probabilities under $\Psi_{1}$: $l_{k}(\mathcal{D}_{c,1,r})=\text{logit}(\tau_{k}(\mathcal{D}_{c,1,r}))$, where the subscript $k$ corresponds to the individual estimand. If a computational method is used to generate posterior samples, we recommend calculating posterior probabilities using a nonparametric kernel density estimate of the posterior distribution so that these logits are finite. We construct linear approximations separately for each estimand using these logits in Line 8. We use these linear approximations to estimate logits of posterior probabilities for new values of $c$ as $\hat{l}_{k}(\mathcal{D}_{c,1})$. We place a hat over the $l$ here to convey that this logit was estimated using a linear approximation instead of a sample with $c$ clusters. This notation also removes the subscript $r$ since the linear approximations are not created based on samples $\mathcal{D}_{c_{0},1,r}$ and $\mathcal{D}_{c_{1},1,r}$ that share the same index $r$.
@@
Given the linear trend in the proxy sampling distribution quantiles discussed in Section 3, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the value of the individual estimand $\delta_{r,k}$ is similar for all $\boldsymbol{\theta}_{r}\sim\Psi_{1}$. When $\Psi_{1}$ is nondegenerate, the process in Line 8 can be modified. We instead split the logits of the posterior probabilities for each $c$ value into subgroups based on the order statistics of their $\delta_{r,k}$ values before constructing the linear approximations. Our linear approximations concern the individual estimands in $\boldsymbol{\delta}(\boldsymbol{\theta})$; however, we must account for the dependence between the individual estimands in the joint sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ when considering new values of $c$. We do so by grouping the estimated logits $\hat{l}_{k}(\mathcal{D}_{c,1})$ across estimands based on the sample $\mathcal{D}_{c_{0},1,r}$ that defined their linear approximation. In Line 10, we find the smallest value of $c$ such that our estimate for $\mathbb{E}_{\Psi_{1}}[\Pr(\nu(\mathcal{D}_{c})=1~{}|~{}\boldsymbol{\theta})]$ based on the indicators $\{\hat{\boldsymbol{\nu}}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ that correspond to $\{\hat{\boldsymbol{\tau}}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ is at least $\Gamma_{1}$.
@@
We did not estimate the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c_{1}})$ under $\Psi_{0}$ in Algorithm 1. In Section 5, we consider models $\Psi_{0}$ that assign all weight to $\boldsymbol{\theta}_{r}$ values such that $\delta_{k}(\boldsymbol{\theta}_{r})$ equals $\delta_{L,k}$ or $\delta_{U,k}$ for all $k=1,\dots,K$. For such models $\Psi_{0}$, large-sample results (Bernardo and Smith, 2009; Golchi and Willard, 2024) guarantee that the probability of making the incorrect decision for a particular set of thresholds $\boldsymbol{\gamma}$ is roughly constant across a range of large $c$ values. If using a more general model $\Psi_{0}$, Algorithm 1 can be adapted to implement the process in Lines 6 to 9 under $\Psi_{0}$ to efficiently estimate the sampling distribution of $\boldsymbol{\tau}(\mathcal{D}_{c})$ for new values of $c$. These estimated sampling distributions could be used to choose an optimal set of thresholds $\boldsymbol{\gamma}$ for each sample size $c$ considered.
@@
Lastly, we quantify the impact of simulation variability on the sample size recommendation by constructing bootstrap confidence intervals for the optimal sample size. We construct these confidence intervals by sampling $m$ times with replacement from $\{\boldsymbol{\tau}(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ and $\{\boldsymbol{\tau}(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$ obtained in Algorithm 1. We note that posterior probabilities for each of the $K$ individual estimands are resampled jointly. We obtain a new sample size recommendation by implementing the process in Lines 7 to 10 of Algorithm 1 with the bootstrap estimates of the sampling distributions at $c_{0}$ and $c_{1}$. This process is repeated $M$ times, and a bootstrap confidence interval for the optimal $c$ is calculated using the percentile method (Efron, 1982). The width of this confidence interval can help inform the choice for the number of simulation repetitions $m$. The suitable coverage properties of such confidence intervals were confirmed in Hagar and Stevens (2025).

### v3

@@ 5 Sample Size Determination Procedure
2:  Compute $\{\tau(\mathcal{D}_{c_{0},0,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{0}$.
3:  Choose $\gamma$ to ensure $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\tau(\mathcal{D}_{c_{0},0,r})\geq\gamma\}\leq\alpha$.
4:  Compute $\{\tau(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{1}$.
5:  If $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\tau(\mathcal{D}_{c_{0},1,r})\geq\gamma\}\geq 1-\beta$, choose $c_{1}<c_{0}$. If not, choose $c_{1}>c_{0}$.
6:  Compute $\{\tau(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$ obtained with $\boldsymbol{\theta}_{r}\sim\Psi_{1}$.
7:  for $r$ in 1:$m$ do
8:   Join the $r^{\text{th}}$ order statistics of $\{l(\mathcal{D}_{c_{0},1,k})\}_{k=1}^{m}$ and $\{l(\mathcal{D}_{c_{1},1,k})\}_{k=1}^{m}$ with a line to obtain $\hat{l}(\mathcal{D}_{c,1,r})$      for new $c$ values.
9:  Obtain $\{\hat{\tau}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ as the inverse logits of the estimates $\{\hat{l}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$.
@@
10:  Find $c_{2}$, the smallest $c\in\mathbb{Z}^{+}$ such that $m^{-1}\sum_{r=1}^{m}\mathbb{I}\{\hat{\tau}(\mathcal{D}_{c,1,r})\geq\gamma\}\geq 1-\beta$.
@@
In Line 3 of Algorithm 1, we choose a suitable threshold $\gamma$ to ensure the estimate for $\mathbb{E}_{\Psi_{0}}[\Pr(\tau(\mathcal{D}_{c})\geq\gamma~|~\boldsymbol{\theta})]$ based on the formula in (7) is at most $\alpha$. In Line 5, we choose a second number of clusters $c_{1}$ at which to estimate the sampling distribution of $\tau(\mathcal{D}_{c})$ under $\Psi_{1}$. This cluster count could be selected using various strategies. For instance, logits $\{l(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ can be estimated for new $c$ values using lines that pass through the points $\{(c_{0},l(\mathcal{D}_{c_{0},1,r}))\}_{r=1}^{m}$ with the limiting slopes from Theorem 1; the value for $c_{1}$ could be found as the smallest cluster count such that power (or assurance) based on these estimated logits is at least $1-\beta$. An alternative strategy is to select $c_{0}$ and $c_{1}$ to facilitate visualization of the operating characteristics over a range of relevant cluster counts.
@@
We emphasize that all posterior probabilities approximated in Lines 2 to 6 of Algorithm 1 are obtained by simulating data according to $\Psi_{0}$ or $\Psi_{1}$. One strength of our methodology is that it can be integrated with any computational or analytical method to estimate posterior probabilities. Lines 7 to 9 compute logits of these probabilities under $\Psi_{1}$: $l(\mathcal{D}_{c,1,r})=\text{logit}(\tau(\mathcal{D}_{c,1,r}))$. If a computational method is used to generate posterior samples, we recommend calculating posterior probabilities using a nonparametric kernel density estimate of the posterior distribution so that these logits are finite. We construct linear approximations using these logits in Line 8. We use these linear approximations to estimate logits of posterior probabilities for new values of $c$ as $\hat{l}(\mathcal{D}_{c,1,r})$. We place a hat over the $l$ here to convey that this logit was estimated using a linear approximation instead of a sample with $c$ clusters.
@@
Given the linear trend in the proxy sampling distribution quantiles discussed in Section 4, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the value of the individual estimand $\delta_{r}$ is similar for all $\boldsymbol{\theta}_{r}\sim\Psi_{1}$. When $\Psi_{1}$ is nondegenerate, the process in Line 8 can be modified. We instead split the logits of the posterior probabilities for each $c$ value into subgroups based on the order statistics of their $\delta_{r}$ values before constructing the linear approximations. In Line 11, we find the smallest value of $c$ such that our estimate for $\mathbb{E}_{\Psi_{1}}[\Pr(\tau(\mathcal{D}_{c})\geq\gamma~|~\boldsymbol{\theta})]$ based on $\{\hat{\tau}(\mathcal{D}_{c,1,r})\}_{r=1}^{m}$ is at least $1-\beta$.
@@
We did not estimate the sampling distribution of $\tau(\mathcal{D}_{c_{1}})$ under $\Psi_{0}$ in Algorithm 1. We generally consider models $\Psi_{0}$ that assign all weight to $\boldsymbol{\theta}_{r}$ values such that $\delta(\boldsymbol{\theta}_{r})$ equals $\delta_{L}$ or $\delta_{U}$. For such models $\Psi_{0}$, large-sample results (Bernardo and Smith, 2009; Golchi and Willard, 2024) guarantee that the expected type I error rate for a particular choice of $\gamma$ is roughly constant across a range of large $c$ values. If using a more general model $\Psi_{0}$, Algorithm 1 can be adapted to implement the process in Lines 6 to 10 under $\Psi_{0}$ to efficiently estimate the sampling distribution of $\tau(\mathcal{D}_{c})$ for new values of $c$. These estimated sampling distributions could be used to choose an optimal threshold $\gamma$ for each cluster count $c$ considered.
@@
Lastly, we quantify the impact of simulation variability on the cluster count recommendation by constructing bootstrap confidence intervals for the optimal cluster count. We construct these confidence intervals by sampling $m$ times with replacement from $\{\tau(\mathcal{D}_{c_{0},1,r})\}_{r=1}^{m}$ and $\{\tau(\mathcal{D}_{c_{1},1,r})\}_{r=1}^{m}$ obtained in Algorithm 1. We obtain a new cluster count recommendation by implementing the process in Lines 7 to 11 of Algorithm 1 with the bootstrap estimates of the sampling distributions at $c_{0}$ and $c_{1}$. This process is repeated $M$ times, and a bootstrap confidence interval for the optimal $c$ is calculated using the percentile method (Efron, 1982). The width of this confidence interval can help inform the choice for the number of simulation repetitions $m$. The suitable coverage properties of such confidence intervals were confirmed in Hagar and Stevens (2025).